# A Spatial-statistical model to analyse historical rutting data

### v2

@@ 4.4.2 Evaluating pavement subgrade support on expected rutting
We examine bearing capacity, surface curvature index (SCI), and base curvature index (BCI) for road segments 2100-2700 using 2021 data, indicative of construction quality and soil pressure sustainability over the past decade (Figure 6, graph (c)). Bearing capacity values exceeding 16 tons (       ) suggest a robust foundation design. The SCI to BCI ratio (Figure 6, graphs (d) and (e)) reveals deflection location within the pavement layers. High rutting may result from factors like side slope height (data unavailable), inadequate drainage, or sub-base layer deformation during thawing periods. Bearing capacity data, sampled and temperature-corrected during summer, may inflate measurements compared to thawing periods. Varying construction material quality along the road stretch (segments 1500-3355) warrants investigation into surface and subsoil layer deflections. An SCI/BCI ratio below 12 (       ) indicates robust construction. Inconsistent ratios suggest layer deformation, with data indicating uneven construction on E14, featuring deeper base layer deflections.
In conclusion, pavement type significantly influences rutting, recommending stronger binder types for extended lifespan.
@@ 5 How road authorities can make use of this model
The model can investigate contributing factors to rutting beyond pavement type, such as drainage issues, heavy traffic, and road design flaws. Comparisons with non-spatial models emphasize the importance of spatial factors in explaining rutting variations, reinforcing the need for spatial models for accurate inference and predictions. The estimated model parameters and credible intervals provide a robust foundation for data-driven decision-making, enabling road authorities to make informed choices regarding maintenance priorities and infrastructure improvements.
@@ 6 Conclusion
This research offers valuable insights into the rutting phenomenon by elucidating both well-established factors and those that remain unexplained. It shows the importance of incorporating spatial information to accurately estimate rutting variability, showing that observations in close proximity exhibit higher similarity than those at a distance. To achieve a comprehensive understanding, it is imperative to consider both observable elements like traffic and temperature, as well as unobservable factors like weather, and integrate spatial rutting components.
Moreover, the study confirms the significant impact of asphalt type selection on rutting performance, with rut-resistant mixtures proving more effective in mitigating rutting. Lane width also influences pavement rutting, with narrower roads experiencing more concentrated traffic loads and, consequently, increased rutting. Conversely, wider roads distribute traffic loads over a larger area, reducing rut formation. The analysis reveals that wider road sections tend to experience higher rut depths compared to narrower road stretches. This result can be attributed to the increased traffic volumes and higher wheel loads present on wider roads, which contribute to more pronounced rutting. To address this issue and minimize rutting problems over time, it is advisable to integrate a life-cycle cost-benefit perspective into road design and planning processes.
Further research is encouraged to explore additional factors influencing rutting, such as bearing capacity, ditch depth for water outflow, and lane width. This would enhance the credibility of research results and the applicability of models to other road networks. Additionally, evaluating the long-term performance of rut-resistant asphalt mixtures under various traffic and environmental conditions would be worthwhile.
@@ Appendix A Inference from model parameters
Table 4 compares estimates from the proposed spatial and non-spatial models. The results show that the spatial models account for uncertainty in factors that vary in space, e.g., weather and other explanatory variables such as bearing capacity and roadside ditches not included in the model. Including more explanatory variables reduces the ‘noise‘ term ($\hat{\sigma}_{\varepsilon}$) and random yearly effect ($\hat{\sigma}_{\gamma}$).
@@ A.2 Marginal Posterior Distributions of Explanatory Variables
Figure A.2 gives the marginal posterior distributions of asphalt type, rut depth from the previous year and lane width. Narrow bands of the distributions indicate that the uncertainty is smaller. Graph (a) shows a clear shift in asphalt type’s contribution to rutting from the spatial and non-spatial models. This is because asphalt type is spatially varying. Stone mastic asphalt (Sma) gave lower rutting values because it has a stronger binder.
(a) Pavement type
@@
Figure A.2: Marginal posterior distributions of pavement type (solid lines are the spatial models, and the dotted lines are the non-spatial models), rut depth from the previous year and lane width.
@@ A.3 Marginal Posterior Distributions of Random Effects
Figure A.3 gives the marginal posterior distributions of the standard deviation of the Gaussian observations ($\hat{\sigma}_{\varepsilon}$), the standard deviation of the random yearly effect ($\hat{\sigma}_{\gamma}$), the range of the common spatial field $\hat{\rho}$, and the standard deviation of the common spatial field $\hat{\sigma}_{\omega(s)}$. The solid lines represent the spatial models and the dotted lines represent the non-spatial models. Narrower densities suggest smaller estimated uncertainty. We find that some of the uncertainty of the Gaussian observations $(\hat{\sigma}_{\varepsilon})$ and random yearly effect $(\hat{\sigma}_{\gamma})$ go into the spatial effect $\{\hat{\sigma}_{\omega(s)}\}$, and this can be interpreted as some of the unobserved variability (e.g., weather, temperature) being explained by the spatial rutting component; see Table 3 and Figure A.3 (a) and (b). This is confirmed by the smaller estimates of the ‘noise’ term $(\hat{\sigma}_{\varepsilon})$ compared with the non-spatial model (Table 4).
We examine the effects of the range of the common spatial field ($\hat{\rho}$), and the standard deviation of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$. As more explanatory variables are added to model (e.g., rut depth), the variability of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$ decreases; see Table 4 and Figure A.3, graph (d).
In Model 2, rut depth from the previous year is included, and this effect is large with the credible interval not covering zero. This indicates that deeper rut depths from the previous year would result in increased rutting; Figure A.2, graph (b). In addition, the spatial range for the common spatial field is, generally, larger for Model 2 compared with Model 3 (the simpler model), where rut depth is included (see Table LABEL:tab:range_for_each_year_appendix in Supplementary Material LABEL:supplementary:addional_results) for estimated $\hat{\rho}_{t}$ for each year). This is so, presumably, because of influenced factors such as varying weather patterns. But including lane width reduces the effect of rut depth on expected rutting.
Refer to caption (a) Marginal posterior distributions of the standard deviation of the Gaussian observations $\hat{\sigma}_{\varepsilon}$. Refer to caption (b) Marginal posterior distributions of the standard deviation of the random yearly effect $\hat{\sigma}_{\gamma}$. Refer to caption (c) Marginal posterior distributions of the nominal range of the common spatial field $\hat{\rho}$ Refer to caption (d) Marginal posterior distributions of the standard deviation of the common spatial field $\hat{\sigma}_{\omega(s)}$.

### v3

@@ 4.4.2 Evaluating pavement subgrade support on expected rutting
We examine bearing capacity, surface curvature index (SCI, micrometre), and base curvature index (BCI) for road segments 42000-54000 using 2021 data, indicative of construction quality and soil pressure sustainability over the past decade. Bearing capacity is a measure of the deflection caused by a rolling load axle on pavement. It is measured using the Rapid Pavement Tester (Raptor, https://www.ramboll.com/rst/rst-products), which continuously collects data on the pavement structure. A value exceeding 16 metric tons (       ) suggests a robust foundation design (Figure 5, graph (c)). The SCI to BCI ratio (Figure 5, graphs (d) and (e)) reveals deflection location within the pavement layers. High rutting may result from factors like side slope height (data unavailable), inadequate drainage, or sub-base layer deformation during thawing periods. Bearing capacity data, sampled and temperature-corrected during summer, may inflate measurements compared to thawing periods. Varying construction material quality along the road stretch (30000-67100 metres) warrants investigation into surface and subsoil layer deflections. An SCI/BCI ratio below 12 (       ) indicates robust construction. Inconsistent ratios suggest layer deformation, with data indicating uneven construction on E14, featuring deeper base layer deflections.
In conclusion, asphalt type significantly influences rutting, recommending more wear resistant and deformation resistant binder types for extended lifespan.
@@ 5 How road authorities can make use of this model
The model can investigate contributing factors to rutting beyond asphalt type, such as drainage issues, heavy traffic, and road design flaws. Comparisons with non-spatial models emphasize the importance of spatial factors in explaining rutting variations, reinforcing the need for spatial models for accurate inference and predictions. The estimated model parameters and credible intervals provide a robust foundation for data-driven decision-making, enabling road authorities to make informed choices regarding maintenance priorities and infrastructure improvements.
@@ 6 Conclusion
This research provides valuable insights into the rutting phenomenon by elucidating both well-established factors and those that remain unexplained. It demonstrates the importance of incorporating spatial information to accurately estimate rutting variability, showing that observations in close proximity exhibit higher similarity than those at a distance. To achieve a comprehensive understanding, it is imperative to consider both observed quantities as well as unobserved quantities. Due to Norway’s specific climatic context, where temperatures rarely exceed 30 ${}^{\circ}\mathrm{C}$ and annual averages remain below 20 ${}^{\circ}\mathrm{C}$, temperature was not included in our rutting model. The maximum average mean temperature per month is less than 18 ${}^{\circ}\mathrm{C}$, indicating that temperature extremes, which could significantly impact rutting by accelerating asphalt softening, are infrequent. Therefore, we focused on more relevant factors such as asphalt type, prior rut depth, and traffic load, which exhibit significant spatial and temporal variability. This approach provides a more accurate and contextually appropriate understanding of the rutting phenomenon in Norway.
The study confirms the significant impact of asphalt type selection on rutting performance, with rut-resistant mixtures proving more effective in mitigating rutting. It also highlights the influence of lane width on pavement rutting. Narrower roads experience more concentrated traffic loads, leading to increased rutting compared to a wider road that is otherwise similar (i.e., same AADT, asphalt type and rut depth levels). However, wider roads, while distributing traffic loads over a larger area, tend to experience higher rut depths, likely due to increased traffic volumes and higher wheel loads.
To address rutting and minimize its effects over time, integrating a life-cycle cost-benefit perspective into road design and planning processes is advisable. Further research should explore additional factors influencing rutting, such as bearing capacity, ditch depth for water outflow, and lane width, to enhance the credibility of research results and model applicability to other road networks. Additionally, evaluating the long-term performance of rut-resistant asphalt mixtures under various traffic and environmental conditions would be beneficial. This comprehensive approach ensures that our model provides more meaningful insights and practical recommendations for mitigating rutting in similar climatic regions.
@@ Appendix A Inference from model parameters
Table 4 compares estimates from the proposed spatial and non-spatial models. The spatial models account for spatially varying uncertainties, such as weather and explanatory variables, e.g., bearing capacity. Including more explanatory variables reduces the noise term ($\hat{\sigma}_{\varepsilon}$) and the random yearly effect ($\hat{\sigma}_{\gamma}$).
@@ A.2 Marginal Posterior Distributions of Explanatory Variables
Figure A.2 gives the marginal posterior distributions of asphalt type, rut depth from the previous year and lane width. Narrow bands of the distributions indicate that the uncertainty is smaller. Graph (a) shows a clear shift in asphalt type’s contribution to rutting from the spatial and non-spatial models. This is because asphalt type is spatially varying. Stone mastic asphalt (Sma) gave lower rutting values because it has a more wear resistant binder.
(a) Asphalt type
@@
Figure A.2: Marginal posterior distributions of asphalt type (solid lines are the spatial models, and the dotted lines are the non-spatial models), rut depth from the previous year and lane width.
@@ A.3 Marginal Posterior Distributions of Random Effects
Figure A.3 presents the marginal posterior distributions of the standard deviation of Gaussian observations ($\hat{\sigma}_{\varepsilon}$), random yearly effect ($\hat{\sigma}_{\gamma}$), range of the common spatial field ($\hat{\rho}$), and standard deviation of the common spatial field ($\hat{\sigma}_{\omega(s)}$). Solid and dotted lines represent spatial and non-spatial models, respectively, with narrower densities indicating smaller estimated uncertainties. The spatial model captures some of the uncertainties in Gaussian observations and yearly effects, suggesting that unobserved variability (e.g., weather) is partly explained by the spatial rutting component. This is evidenced by smaller noise term estimates ($\hat{\sigma}_{\varepsilon}$) in the spatial model (Table 4).
As more explanatory variables are added (e.g., rut depth), the variability of the spatial field ($\hat{\sigma}_{\omega(s)}$) decreases (Figure A.3, graph (d)). In Model 2, including previous year’s rut depth gives a large effect, indicating deeper ruts lead to increased rutting (Figure A.2, graph (b)). The spatial range ($\hat{\rho}$) is generally larger in Model 2 compared to Model 3, likely due to varying weather patterns, but including lane width reduces the effect of rut depth on expected rutting.
(a) Marginal posterior distributions of the standard deviation of the Gaussian observations $\hat{\sigma}_{\varepsilon}$.
(b) Marginal posterior distributions of the standard deviation of the random yearly effect $\hat{\sigma}_{\gamma}$.
(c) Marginal posterior distributions of the nominal range of the common spatial field $\hat{\rho}$
(d) Marginal posterior distributions of the standard deviation of the common spatial field $\hat{\sigma}_{\omega(s)}$.
@@ A.4 Area with high rutting rates
A.5 Average mean air surface temperature
In Norway, the temperature rarely exceeds 30 ${}^{\circ}\mathrm{C}$, and on average, monthly temperatures do not exceed 20 ${}^{\circ}\mathrm{C}$. This limited variation in high temperatures suggests that temperature may not play a critical role in the rutting process within this specific climatic context. The maximum average mean temperature per month or over years in Norway is less than 18 ${}^{\circ}\mathrm{C}$ (and less than $13.5\ ^{\circ}\mathrm{C}$ for all years, except 2104 and 2018). This indicates that temperature extremes, which could significantly influence rutting, are infrequent.
Figure A.5: Average mean air surface temperature (${}^{\circ}\mathrm{C}$,  ) for years 2011-2020. The black circles ($\bullet$) gives the average mean monthly air surface temperature and the coloured dots indicated in the legend gives the mean monthly for each year. The dashed black line gives a temperature value of $13.5^{\circ}\mathrm{C}$ (see https://climateknowledgeportal.worldbank.org/country/norway/trends-variability-historical for further details).